find value of x

log 2 base of x+log(x-7)=3 is the same as logx(2) + log(x-7) = 3

the graph is shown below:



Answer by Theo(13342)

log2(x) + log(x-7) = 3 doesn't have a solution as far as i can see.

i couldn't solve it algebraically, but i could solve it graphically.

the solution is x = 794.40479.

when x = 794.40479, the graph of y = logx(2) + log(x-7) is equal to 3.

it will then intersect with the line of y = 3.

logx(2) is the same as log(2) / log(x), so the equation that was graphed was y = log(2)/log(x) + log(x-7)

log without a base is assumed to be a base of 10 by convention.

here's the graph.

$$$